# GQE: Generalized Query Expansion for Enhanced Text-Video Retrieval

### v1

@@ 3.3 GQE: Query Expansion of Testing
Query Selection. Before delving into query selection, we first explore the potential benefits of incorporating the generated queries by calculating an upper bound performance. In particular, from the set of $\{Q,Q_{1}^{test},...,Q_{n}^{test}\}$, we select the oracle query that yields the best retrieval result. We found that a huge improvement can be achieved via the oracle query selection. In MSR-VTT dataset, the Rank-1 metric is boosted from $46.7\%$ to $61.4\%$ (see Tab. 6). This surprising result reveals that: i) query expansion has significant potential for improving text-video retrieval; ii) selecting an effective query can simultaneously reduce computational costs and enhance retrieval performance.
@@
The result motivates us to investigate effective query selection. When defining effective query selection, our primary considerations are two aspects: relevance and diversity. The former aspect demands that expanded queries should accurately convey semantic meaning with minimal illusions. The latter emphasizes that the text queries should be expressed in a variety of ways, such that more novel views can be included. Often, these two aspects exhibit some degree of contradiction. Relevance may limit diversity, whereas increased diversity also increases the risk of introducing illusions. In this work, we firstly control the relevance within a certain range by designing an appropriate prompt, which forces that LLM query generation strictly adheres to the factual information in the original query. As shown in Fig. 3, the output queries are constrained within a range of relevance by doing this. After that, we sample a certain number of queries from this pool. The goal of this step is to maximize the diversity within this range. In the process, we drew inspiration from point cloud research work [45] and transformed it into the Farthest Query Sampling (FQS) algorithm on the query feature space.
Figure 3: Illustration of Farthest Query Sampling (FQS) algorithm. The queries are distributed within a range of relevance. The blue point (original query) is set as the first query. In this example, FQS sample 2 expanded queries to enrich diversity by considering the distance in the embedding space.
@@
The main idea is to iteratively select queries in a way that maximizes the minimum distance between the selected queries. The algorithm starts with an initial query and repeatedly selects the query that is farthest from all the previously selected queries. This process continues until the desired number of queries is selected. Note that the result of this algorithm is affected by the initial query. We set the original user query as the initial query to stabilize the algorithm and provide double insurance for relevance. The formal expression is Eq. 9, where $Q$ is the original query, $\{Q_{1}^{*},...,Q_{k}^{*}\}$ are the expanded queries, $k$ is the number of query selection ($k\leq n$). The distance is computed at an embedding space.
@@ 4 Experiments
Setups. We employ MSR-VTT [66], MSVD [64], LSMDC [49], and VATEX [61] to evaluate our method. We use standard retrieval metrics: recall at rank K (R@K, higher is better), median rank (MdR, lower is better), and mean rank (MnR, lower is better) to evaluate the performance. The model backbone is initialized from pre-trained CLIP ViT-B/32 and undergo end-to-end training on each dataset. More details of experiment setup and implementation are elaborated in the appendix.
@@ 4.1 Comparison with the State-of-the-Art
Table 1: Text-to-video retrieval results on the 1K-A test set of MSR-VTT 1K [65]. *Denotes that the method uses DSL [12] post-processing.
@@
CLIP4Clip [38] Neurocomp.’22 44.5 71.4 81.6 2.0 15.3
CAMoE* [12] arXiv’21 47.3 74.2 84.5 2.0 11.9
CLIP2Video [17] arXiv’21 45.6 72.6 81.7 2.0 14.6
X-Pool [23] CVPR’22 46.9 72.8 82.2 2.0 14.3
QB-Norm [7] CVPR’22 47.2 73.0 83.0 2.0 -
TS2-Net [36] ECCV’22 47.0 74.5 83.8 2.0 13.0
DRL [58] arXiv’22 47.4 74.6 83.8 2.0 -
UATVR [16] ICCV’23 47.5 73.9 83.5 2.0 12.9
Cap4Video [62] CVPR’23 49.3 74.3 83.8 2.0 12.0
TEFAL [24] ICCV’23 49.4 75.9 83.9 2.0 12.0
CLIP-VIP [68] ICLR’23 50.1 74.8 84.6 1.0 -
GQE 52.1 76.8 86.3 1.0 9.7
GQE* 54.1 81.6 89.7 1.0 8.0
CLIP-VIP [68]+GQE 53.2 79.3 88.2 1.0 10.2
CLIP-VIP [68]+GQE* 58.9 83.4 89.5 1.0 8.3
@@
CLIP2TV [22] arXiv’21 48.3 74.6 82.8 2.0 14.9
CenterCLIP [71] SIGIR’22 48.4 73.8 82.0 2.0 13.8
TS2-Net [36] ECCV’22 49.4 75.6 85.3 2.0 13.5
DRL [58] arXiv’22 50.2 76.5 84.7 1.0 -
UATVR [16] ICCV’23 50.8 76.3 85.5 1.0 12.4
Cap4Video [62] CVPR’23 51.4 75.7 83.9 1.0 12.4
CLIP-VIP [68] ICLR’23 54.2 77.2 84.8 - -
GQE 54.6 79.5 86.7 1.0 9.1
GQE* 57.1 81.9 89.2 1.0 8.0
Table 2: Comparisons with SOTA of t2v retrieval on MSVD [63].
@@
CE [35] 19.8 49.0 63.8 6.0 -
SUPPORT [42] 28.4 60.0 72.9 4.0 -
CLIP [47] 37.0 64.1 73.8 3.0 -
Frozen [5] 33.7 64.7 76.3 3.0 -
TMVM [33] 36.7 67.4 81.3 2.5 -
CLIP4Clip [38] 45.2 75.5 84.3 2.0 10.3
UATVR [16] 46.0 76.3 85.1 2.0 10.4
X-Pool [23] 47.2 77.4 86.0 2.0 9.3
DRL [58] 48.3 79.1 87.3 2.0 -
GQE 51.1 81.1 88.6 1.0 7.2
Table 3: Comparisons with SOTA of t2v retrieval on VATEX [61].
@@
HGR [11] 35.1 73.5 83.5 2.0 -
CLIP [47] 39.7 72.3 82.2 2.0 12.8
SUPPORT [42] 44.9 82.1 89.7 1.0 -
CLIP4Clip [38] 55.9 89.2 95.0 1.0 3.9
Clip2Video [17] 57.3 90.0 95.5 1.0 3.6
QB-Norm [7] 58.8 88.3 93.8 1.0 -
TS2-Net [36] 59.1 90.0 95.2 1.0 3.5
TEFAL [24] 61.0 90.4 95.3 1.0 3.8
UATVR [16] 61.3 91.0 95.6 1.0 3.3
GQE 63.3 92.4 96.8 1.0 2.8
Table 4: Comparisons with SOTA of t2v retrieval on LSMDC [50].
@@
CE [35] 11.2 26.9 34.8 25.3 -
MMT [21] 12.9 29.9 40.1 19.3 75.0
NoiseE [1] 6.4 19.8 28.4 39.0 -
Straight-CLIP [43] 11.3 22.7 29.2 56.5 -
MDMMT [15] 18.8 38.5 47.9 12.3 58.0
Frozen [5] 15.0 30.8 39.8 20.0 -
TeachText-CE+ [13] 17.2 36.5 46.3 13.7 -
CLIP4Clip [38] 22.6 41.0 49.1 11.0 61.0
XPool [23] 25.2 43.7 53.5 8.0 53.2
CLIP-VIP [68] 25.6 45.3 54.4 8.0 -
TEFAL [24] 26.8 46.1 56.5 7.0 44.4
GQE 28.3 49.2 59.1 6.0 37.1
In this section, we compare our Generalized Query Expansion (GQE) method with recent state-of-the-art methods on four benchamarks. Tab. 1 shows the comparisons on MSR-VTT. Our method sets a new state-of-the-art performance in text-to-video retrieval for both ViT-B/32 and ViT-B/16 backbones, surpassing previous methods significantly. For instance, we achieve a remarkable +7.6% higher R@1 compared to CLIP4Clip when using the same ViT-B/32 backbone for text-to-video retrieval, without any post-processing techniques. In comparison to methods that also utilize auxiliary captions, such as Cap4Video [62] and CLIP-VIP [68], our approach consistently achieves higher performance with both ViT-B/32 and ViT-B/16 backbones. It’s worth noting that CLIP-VIP uses the large-scale video-text dataset HowTo100M [39] for pre-training. Despite this, our method still outperforms it, providing further evidence of effectiveness of our approach. We further show that, as a data-centric method, GQE is compatible with stronger models, such as CLIP-VIP [68], and post-processing techniques, such as DSL [12], yielding significantly higher performance.
Furthermore, we conduct evaluations on other benchmarks, including MSVD (Tab. 3), VATEX (Tab. 3), and LSMDC (Tab. 4). To ensure a fair comparison, all experiments in these three datasets are conducted using the ViT-B/32 backbone. GQE consistently improves performance across different datasets, ultimately achieving new state-of-the-art results. Overall, our method’s consistent performance improvements across these four benchmarks demonstrate the effectiveness of GQE.
@@ 4.2 Ablation Study
Table 5: Ablation study on query expansion of training set on MSR-VTT [65]. We compare our GQE (Rows 4-6) with various training query expansion settings.
@@
2 Global Video Captions 46.9 73.0 81.9 2.0 13.9
3 Fixed Time Interval 47.0 72.2 82.3 2.0 14.2
4 GQE (Training) 47.7 71.9 82.1 2.0 13.9
5 GQE (Testing) 50.2 78.0 86.9 1.0 10.1
6 GQE (Full) 52.1 76.8 86.3 1.0 9.7
Table 6: Ablation study on query generation and aggregation. We compare various strategies of aggregating the results of testing query expansion, under two query generation solutions.
@@
4.2.1 Query Expansion of Training Set
In Tab. 6, we study the impact of the query expansion of training set. Row 1 is the baseline model trained on the original MSR-VTT [65], without training query expansion. Row 2 shows results with the model pre-trained on video captions provided by a previous work [62], which uses a zero-shot video captioning model ZeroCap [53]. In Row 3, we design a strawman setting, in which we uniformly break a video in fixed intervals and do captioning for each chunk. The text-video retrieval model is then pre-trained on this expanded caption set. In Rows 4 and Row 6, the model is pre-trained on comprehensive video captions generated by our approach. To ensure that the quantity of data does not lead to unfair comparison, we generate a comparable number of captions for Rows 2 to 6, resulting in approximately 45K additional queries for the 9K training videos in the MSR-VTT dataset. Comparing Rows 1 to 4 suggests that incorporating extra text queries in the training data is generally beneficial. Particularly, GQE (Training) is shown to be more effective, demonstrating a 1.0% performance increase. We attribute this improvement to the proposed approach, which effectively reduces the information imbalance gap by covering more video scenes. When comparing GQE (Testing) and GQE (Full), we observe that GQE (Training) is not only effective on the baseline model but also enhances performance based on GQE (Testing), resulting in a 1.9% increase. These observations confirm the effectiveness of training set query expansion.
Table 7: Ablation study on query selection, including the selection algorithm and the choice of parameter $k$. We report only R@1 metric for simplicity.

### v2

@@ 3.3 Query Selection Mechanism
We first introduce the concept of “Oracle Query". From the set of $\{Q,Q_{1}^{test},...,Q_{n}^{test}\}$, we always select the one that yields the best retrieval performance. Since this process requires access to the ground truth video-text matching, which is impractical in application, we only use this approach to demonstrate the potential of text query enrichment and query selection. It can be regarded as an upper bound performance. We found that this oracle query selection can achieve dramatic performance improvement even on a baseline model. In MSR-VTT dataset, by simply replacing the original query with oracle query, the Rank-1 metric is boosted from $46.7\%$ to $61.4\%$ (see Tab. 6). This surprising result reveals two key insight: i) text enrichment in text phase has significant potential for improving text-video retrieval; ii) selecting an effective query can simultaneously reduce computational costs and enhance retrieval performance.
Figure 3: Illustration of Farthest Query Sampling (FQS) algorithm. The queries are distributed within a certain range of relevance. The blue point (user query) is set as the root query. At each step, FQS samples the query that is farthest from all previous sampled queries.
@@
The result motivates us to investigate effective query selection. When defining effective query selection, our primary considerations are two aspects: relevance and diversity. The former aspect demands that enriched queries should accurately convey semantic meaning with minimal illusions. The latter emphasizes that the text queries should be expressed in a variety of ways, such that more novel views can be included. Often, these two aspects exhibit some degree of contradiction. Relevance may limit diversity, whereas increased diversity also increases the risk of introducing illusions. In this work, the relevance is controlled in two folds. First, by designing an appropriate prompt, we enforce the LLM to generate enriched text queries strictly adheres to the factual information in the original query, limiting the generated ones in a certain range. Secondly, in the following query selection process, we always primarily include the user query as the initial one, serving as the most relevant and reliable anchor. As shown in Fig. 3, the output queries are constrained within a range of relevance. After that, we sample a certain number of queries from this pool. The goal of this step is to maximize the diversity within this range. In the process, we design a novel Farthest Query Sampling (FQS) algorithm inspired by point sampling in point cloud research Qi et al. (2017).
@@
The main idea is to iteratively select queries in a way that maximizes the minimum distance between the selected queries. The algorithm starts with an initial query and repeatedly selects the query that is farthest from all the previously selected queries. This process continues until the desired number of queries is selected. Note that the result of this algorithm is affected by the initial query. We set the original user query as the initial query provided by the human user to stabilize the algorithm and provide double insurance for relevance. The formal expression is Eq. 9, where $Q$ is the original query, $\{Q_{1}^{*},...,Q_{k}^{*}\}$ are the enriched queries, $k$ is the number of query selection ($k\leq n$). The distance is computed at an embedding space.
@@ 4 Experiments
Setups. We employ MSR-VTT Xu et al. (2016a), MSVD Wu et al. (2017a), LSMDC Rohrbach et al. (2015a), and VATEX Wang et al. (2019) to evaluate our method. We use standard retrieval metrics: recall at rank K (R@K, higher is better), median rank (MdR, lower is better), and mean rank (MnR, lower is better) to evaluate the performance. The model backbone is initialized from pre-trained CLIP ViT-B/32 and undergo end-to-end training on each dataset. More details of experiment setup and implementation are elaborated in the appendix.
@@ 4.1 Comparison with the State-of-the-Art
Table 1: Text-to-video retrieval results on the 1K-A test set of MSR-VTT 1K Xu et al. (2016b). *Denotes that the method uses DSL Cheng et al. (2021) post-processing.
@@
CLIP4Clip Luo et al. (2022) Neurocomp.’22 44.5 71.4 81.6 2.0 15.3
X-Pool Gorti et al. (2022) CVPR’22 46.9 72.8 82.2 2.0 14.3
QB-Norm Bogolin et al. (2022) CVPR’22 47.2 73.0 83.0 2.0 -
TS2-Net Liu et al. (2022) ECCV’22 47.0 74.5 83.8 2.0 13.0
DRL Wang et al. (2022) arXiv’22 47.4 74.6 83.8 2.0 -
UATVR Fang et al. (2023) ICCV’23 47.5 73.9 83.5 2.0 12.9
Cap4Video Wu et al. (2023b) CVPR’23 49.3 74.3 83.8 2.0 12.0
TEFAL Ibrahimi et al. (2023) ICCV’23 49.4 75.9 83.9 2.0 12.0
ProST Li et al. (2023c) ICCV’23 48.2 74.6 83.4 2.0 12.4
UCOFIA Wang et al. (2023c) ICCV’23 49.4 72.1 - - 12.9
UMT Li et al. (2023b) ICCV’23 51.0 76.5 84.2 - -
CLIP-VIP Xue et al. (2022) ICLR’23 50.1 74.8 84.6 1.0 -
T-MASS Wang et al. (2024a) CVPR’24 50.2 75.3 85.1 1.0 11.9
TeachCLIP Tian et al. (2024) CVPR’24 46.8 74.3 - - -
KDProR Zhuang et al. (2024) ECCV’24 49.6 75.1 84.4 2.0 11.6
DITS Wang et al. NeurIPS’24 51.9 75.7 84.6 1.0 11.6
X-Pool Gorti et al. (2022) + Ours 52.1 76.8 86.3 1.0 9.7
X-Pool Gorti et al. (2022) + Ours* 54.1 81.6 89.7 1.0 8.0
CLIP-VIP Xue et al. (2022) + Ours 53.2 79.3 88.2 1.0 10.2
CLIP-VIP Xue et al. (2022) + Ours* 58.9 83.4 89.5 1.0 8.3
@@
CLIP2TV Gao et al. (2021) arXiv’21 48.3 74.6 82.8 2.0 14.9
CenterCLIP Zhao et al. (2022) SIGIR’22 48.4 73.8 82.0 2.0 13.8
TS2-Net Liu et al. (2022) ECCV’22 49.4 75.6 85.3 2.0 13.5
DRL Wang et al. (2022) arXiv’22 50.2 76.5 84.7 1.0 -
UATVR Fang et al. (2023) ICCV’23 50.8 76.3 85.5 1.0 12.4
Cap4Video Wu et al. (2023b) CVPR’23 51.4 75.7 83.9 1.0 12.4
CLIP-VIP Xue et al. (2022) ICLR’23 54.2 77.2 84.8 - -
T-MASS Wang et al. (2024a) CVPR’24 52.7 77.1 85.6 1.0 10.5
X-Pool Gorti et al. (2022) + Ours 54.6 79.5 86.7 1.0 9.1
X-Pool Gorti et al. (2022) + Ours* 57.1 81.9 89.2 1.0 8.0
Table 2: Comparisons with SOTA of t2v retrieval on MSVD Wu et al. (2017b). Grey text denotes using ViT-B/16 backbone.
@@
CE Liu et al. (2019) 19.8 49.0 63.8 6.0
SUPPORT Patrick et al. (2021) 28.4 60.0 72.9 4.0
CLIP Radford et al. (2021) 37.0 64.1 73.8 3.0
Frozen Bain et al. (2021) 33.7 64.7 76.3 3.0
TMVM Lin et al. (2022) 36.7 67.4 81.3 2.5
CLIP4Clip Luo et al. (2022) 45.2 75.5 84.3 2.0
UATVR Fang et al. (2023) 46.0 76.3 85.1 2.0
X-Pool Gorti et al. (2022) 47.2 77.4 86.0 2.0
DRL Wang et al. (2022) 48.3 79.1 87.3 2.0
UCOFIA Wang et al. (2023c) 47.4 77.6 - -
TeachCLIP Tian et al. (2024) 47.4 77.3 - -
Cap4Video Wu et al. (2023b) 51.8 80.8 88.3 1.0
X-Pool Gorti et al. (2022) + Ours 51.1 81.1 88.6 1.0
Table 3: Comparisons with SOTA of t2v retrieval on VATEX Wang et al. (2019). Grey text denotes using ViT-B/16 backbone.
@@
HGR Chen et al. (2020) 35.1 73.5 83.5 2.0
CLIP Radford et al. (2021) 39.7 72.3 82.2 2.0
SUPPORT Patrick et al. (2021) 44.9 82.1 89.7 1.0
CLIP4Clip Luo et al. (2022) 55.9 89.2 95.0 1.0
QB-Norm Bogolin et al. (2022) 58.8 88.3 93.8 1.0
TS2-Net Liu et al. (2022) 59.1 90.0 95.2 1.0
TEFAL Ibrahimi et al. (2023) 61.0 90.4 95.3 1.0
ProST Li et al. (2023c) 60.6 90.5 95.4 1.0
UATVR Fang et al. (2023) 61.3 91.0 95.6 1.0
UCOFIA Wang et al. (2023c) 61.1 90.5 - -
T-MASS Wang et al. (2024a) 63.0 92.3 96.4 1.0
Cap4Video Wu et al. (2023b) 66.6 93.1 97.0 1.0
X-Pool Gorti et al. (2022) + Ours 63.3 92.4 96.8 1.0
Table 4: Comparisons with SOTA of t2v retrieval on LSMDC Rohrbach et al. (2015b).
@@
CE Liu et al. (2019) 11.2 26.9 34.8 25.3 -
MMT Gabeur et al. (2020) 12.9 29.9 40.1 19.3 75.0
NoiseE Amrani et al. (2021) 6.4 19.8 28.4 39.0 -
Straight-CLIP Portillo-Quintero et al. (2021) 11.3 22.7 29.2 56.5 -
MDMMT Dzabraev et al. (2021) 18.8 38.5 47.9 12.3 58.0
Frozen Bain et al. (2021) 15.0 30.8 39.8 20.0 -
TeachText-CE+ Croitoru et al. (2021) 17.2 36.5 46.3 13.7 -
CLIP4Clip Luo et al. (2022) 22.6 41.0 49.1 11.0 61.0
XPool Gorti et al. (2022) 25.2 43.7 53.5 8.0 53.2
CLIP-VIP Xue et al. (2022) 25.6 45.3 54.4 8.0 -
ProST Li et al. (2023c) 24.1 42.5 51.6 9.0
TEFAL Ibrahimi et al. (2023) 26.8 46.1 56.5 7.0 44.4
X-Pool Gorti et al. (2022) + Ours 28.3 49.2 59.1 6.0 37.1
In this section, we compare our method with recent state-of-the-art methods on four benchamarks. Tab. 1 shows the comparisons on MSR-VTT. Our method sets a new state-of-the-art performance in text-to-video retrieval for both ViT-B/32 and ViT-B/16 backbones, surpassing previous methods significantly. For instance, we achieve a remarkable +7.6% higher R@1 compared to CLIP4Clip when using the same ViT-B/32 backbone for text-to-video retrieval, without any post-processing techniques. In comparison to methods that also utilize synthetic text captions, such as Cap4Video Wu et al. (2023b) and CLIP-VIP Xue et al. (2022), our approach consistently achieves higher performance with both ViT-B/32 and ViT-B/16 backbones. It’s worth noting that CLIP-VIP uses the large-scale video-text dataset HowTo100M Miech et al. (2019) for pre-training. Despite this, our method still outperforms it, providing further evidence of effectiveness of our approach. We further show that, as a data-centric method, our method is compatible with stronger models, such as CLIP-VIP Xue et al. (2022), and post-processing techniques, such as DSL Cheng et al. (2021), yielding significantly higher performance.
Furthermore, we conduct evaluations on other benchmarks, including MSVD (Tab. 3), VATEX (Tab. 3), and LSMDC (Tab. 4). To ensure a fair comparison, all experiments in these three datasets are conducted using the ViT-B/32 backbone. Our method consistently improves performance across different datasets, ultimately achieving new state-of-the-art results. Overall, our method’s consistent performance improvements across these four benchmarks demonstrate the its effectiveness.
@@ 4.2 Ablation Study
Table 5: Ablation study on text enrichment of training phase on MSR-VTT Xu et al. (2016b). We compare our method with other captioning-based text enrichment strategies.
@@
2 Global Video Caption 46.9 73.0 81.9 2.0 13.9
3 Fixed Time Interval Caption 47.0 72.2 82.3 2.0 14.2
4 Ours (Training Enrich. Only) 47.7 71.9 82.1 2.0 13.9
5 Ours (Retrieval Enrich. Only) 50.2 78.0 86.9 1.0 10.1
6 Ours (Unified Enrichment) 52.1 76.8 86.3 1.0 9.7
Table 6: Ablation study on text enrichment of the retrieval phase, including query generation and result aggregation strategies.
@@
4.2.1 Text Enrichment of Training Phase
In Tab. 6, we study the impact of text enrichment of training phase. Row 1 is the baseline model trained on the original MSR-VTT Xu et al. (2016b), without query enrichment in training. Row 2 shows results with the model pre-trained on video captions provided by a previous work Wu et al. (2023b), which uses a zero-shot video captioning model ZeroCap Tewel et al. (2022). In Row 3, we design a strawman setting, in which we uniformly break a video in fixed intervals and do captioning for each chunk. The text-video retrieval model is then pre-trained on this enriched caption set. In Rows 4 and Row 6, the model is pre-trained on comprehensive, event-aware video captions generated by our approach. To ensure that the scale of data does not lead to unfair comparison, we generate a comparable number of captions for Rows 2 to 6, resulting in approximately 45K additional queries for the 9K training videos in the MSR-VTT dataset. Comparing Rows 1 to 4 suggests that incorporating extra text queries in the training data is generally beneficial. Particularly, Our (Training Enrich. Only) is shown to be more effective, demonstrating a 1.0% performance increase. We attribute this improvement to the proposed approach, which effectively reduces the information gap by covering more video scenes. To further validate effectiveness, we compare Row 5 Ours (Testing Enrich. Only) and Row 6 Ours (Unified Enrichment). The 1.9% performance gap demonstrate the text enrichment of training again.
Table 7: Ablation study on query selection, including the selection algorithm and the choice of parameter $k$, i.e., number of enriched text queries. We report only R@1 metric for simplicity. The bold number denotes the highest number for each row.